What is the actual elevation angle of the Sun above the horizontal?

An underwater scuba diver sees the Sun at an apparent angle of 45.5° above the horizontal. What is the actual elevation angle of the Sun above the horizontal? (Use1.333 for the index of refraction of water.)
...........° above the horizon
Answer #1

n1sinθ1=n2sinθ2

=>1sinθ1=1.333sin45.5

sinθ1=0.950

=>θ1=sin-1(0.950)

=>θ1=71.945

=>Angle of elevation=90-71.945

=18.055 above the horizontal

answered by: Derric
Answer #2
sint=1.33*sin45.5
t=71.55
answered by: Joelle
Answer #3
sin(90 - x) = 1.33 sin(90-45.5)
answered by: Quay
Answer #4
refraction angle = 90 - 45.5 = 44.5 degrees
sin(90 - x) = 1.33 sin(44.5)
so x = 21.2 degrees
answered by: Amos